OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty} \frac{ x^4*\sin \frac{ 1 }{ x }+x^2 }{ 1+\left| x \right|^3}\]

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

@mathslover ?

mathslover (mathslover):

Yep.. working on it. Did you try using (a^3 + b^3) identity? not sure whether that will work or not.

OpenStudy (anonymous):

x=-1/h and h->0

myininaya (myininaya):

on that |x| is that a cube? I can't read for some reason. I can click on the code.

OpenStudy (anonymous):

yes it is a cube

mathslover (mathslover):

I suggest you simplify the denominator... use the identity.

OpenStudy (anonymous):

that mod x will open with minus?

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

may be divide numerator and denominator by x^3

myininaya (myininaya):

so since x<0 then |x|=-x and since we have |x|^3 then we could replace |x|^3 with (-x)^3=-x^3 now recall that if u->0 then sin(u)/u->1 see if you can use that here divide both top and bottom by 1/x

myininaya (myininaya):

the some l'hopital could be used :)

OpenStudy (zarkon):

for large x \[\Large \frac{ x^4\sin \frac{ 1 }{ x }+x^2 }{ 1+\left| x \right|^3}\approx \frac{ x^4\sin \frac{ 1 }{ x }+x^2 }{ \left| x \right|^3}\] \[\Large=\frac{ x^4\sin \frac{ 1 }{ x }+x^2 }{ x^2\left| x \right|}\] \[\Large=\frac{ x^2\sin \frac{ 1 }{ x }+1 }{ \left| x \right|}\] \[\Large\approx\frac{ x^2\sin \frac{ 1 }{ x } }{ \left| x \right|}=\frac{x}{|x|}x\sin\frac{1}{x}\]

OpenStudy (anonymous):

we have to write x=-1/h and then h->0

OpenStudy (zarkon):

as \(x\to \infty\) you then get \((-1)\cdot 1=-1\)

OpenStudy (zarkon):

x to -infinity you get what i have above

OpenStudy (anonymous):

@Zarkon how did u use the approximation in the first step?

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

@mathslover ?

OpenStudy (zarkon):

in my first step I got rid of the 1 since for large x the 1 really contributes almost nothing

OpenStudy (zarkon):

and by large x I mean large negative (obviously)

OpenStudy (anonymous):

@mathslover ?

myininaya (myininaya):

I would have done it like this: (but this is because i'm not commander data like zarkon (who knows all because he is a superior being to a human) \[\lim_{x \rightarrow - \infty}\frac{\frac{x^4 \sin(\frac{1}{x})}{\frac{1}{x}}+\frac{x^2}{\frac{1}{x}}}{\frac{1}{\frac{1}{x}}-\frac{x^3}{\frac{1}{x}}}\] then use the fact that if u->0 then sin(u)/u->1 it should be pretty easy after this point though.

OpenStudy (sidsiddhartha):

as xtends to - infinity mod(x)=-x now try to put \[x=\frac{ -1 }{ t}\]

OpenStudy (anonymous):

ya @sidsiddhartha

OpenStudy (anonymous):

i was thinking that only

myininaya (myininaya):

or i guess you could have just said \[\lim_{x \rightarrow -\infty}\frac{x^4 \sin(\frac{1}{x})+x^2}{1-x^3}=\lim_{x \rightarrow -\infty}\frac{x^3 \frac{\sin(\frac{1}{x})}{\frac{1}{x}}+x^2}{1-x^3}\]

OpenStudy (sidsiddhartha):